kindly sketch the graph of 2y=x=1 and show me the solution... thanks
Answer by Edwin McCravy(20077)

You typed two equal signs 2y=x=1.  So you probably 
meant to type one of these:

2y+x=1
2y-x=1
2y=x+1
2y=x-1

I will assume it was the first one.  Also you didn't say whether to
use the slope, rise over run, or the point method.  So I'll assume
it was the point method.

We find three (or more) points 

Arbitrarily pick y=0
  2y+x=1
2(0)+x=1
   0+x=1
     x=1, so the line goes through the point (x,y) = (1,0)

Arbitrarily pick x=5
  2y+x=1
2y+(5)=1
  2y+5=1
    2y=-4
     y=-2, so the line goes through the point (x,y) = (5,-2)

Arbitrarily pick y=3
  2y+x=1
2(3)+x=1
   6+x=1
     x=-5, so the line goes through the point (x,y) = (-5,3)


Now we get some graph paper, draw an x-axis and a y-axis.
Then we plot those 3 points:



Now we get a ruler and draw a line through those points:



Really you need only two points, but we always find a third point as
a check, because if they don't line up, we would know we had made a 
mistake.

Edwin